A capacitor C1 is connected to a battery , the battery is removed another capacitor C2 is connected in parallel, (c1!=c2) wat is the charge q?


1
Expert's answer
2021-03-09T15:27:28-0500

Let the battery voltage be "V_0". Then the amount of charge stored by the first capacitor is:


"q_0 = C_1V_0"

After connecting the second capacitor, the total charge remains the same (according to the charge conservation law):


"q = q_0"

The total capacitance of the system now is (parallel connection):


"C = C_1 + C_2"

Thus, the total voltage is:


"V = \dfrac{q}{C} = \dfrac{q_0}{C_1 + C_2} = \dfrac{C_1}{C_1 + C_2}V_0"

Then the charge in the second capacitor is:


"q_2 = C_2V = \dfrac{C_1C_2}{C_1 + C_2}V_0"

Answer. "\dfrac{C_1C_2}{C_1 + C_2}V_0".